A capacitor has a total charge of 6 µC stored. If it were possible to transfer this charge uniformly through a light with a current of 10 mA, how long would the light stay lit?
1
Expert's answer
2011-02-07T08:23:37-0500
From the definition of the electric current, I = Q/t:
t = Q/I = 6x10-6[C]/(10x10-3[A]) = 0.6x10-3 s = 0.6 ms.